OpenStudy (anonymous):

The diagonal of a square is 8 times the square root of 2 units. Which is the perimeter of the square? Answer 8 16 24 32

OpenStudy (anonymous):

hi do you know the formula for the diagonal and perimeter of square?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let the side of a square =s, it has 4 sides therefore perimeter of square p=4s and the diagonal of a square d=sqrt(s^2+s^2)=sqrt(2s^2)=s(sqrt2)=1.4141*s approximately now if given d=8(sqrt2)=s(sqrt2)..we get s=8 now p=4s p=4(8)=32 ans

OpenStudy (anonymous):

any question chris?

OpenStudy (anonymous):

no but thank you